A merchant wants to mix the peanuts with the cashews shown in the illustration to get 48 pounds of mixed nuts to sell at $4.00 per pound. How many pounds of each should the merchant use. Peanuts $3.00/lb. Cashews $6.00/lb.
Found 2 solutions by jorel1380, Alan3354:
Answer by jorel1380(3719)

Let n be the amount of cashews in the mix. Then peanuts would be 48-n. So:
6n+3(48-n)=48(4)
3n=48
n=16
The merchant needs 16 lbs. of cashews and 32 lbs of peanuts for his mix. ☺☺☺☺

Answer by Alan3354(69443)

A merchant wants to mix the peanuts with the cashews shown in the illustration to get 48 pounds of mixed nuts to sell at $4.00 per pound. How many pounds of each should the merchant use. Peanuts $3.00/lb. Cashews $6.00/lb.
----
P + C = 48
3P + 6C = 4*48